OpenStudy (anonymous):

What does it mean to say that the forces acting on an object are balanced? A. The forces all have the same magnitude. B. All the forces combine to give zero net force. C. The sum of the horizontal forces is zero. D. The forces occur as pairs of forces with equal magnitude and direction. **Not too sure, but think it's either C or D? Thank you!!

OpenStudy (danjs):

B

OpenStudy (danjs):

Lets look at each one

OpenStudy (anonymous):

ohh okay:)

OpenStudy (danjs):

When you say the forces are balanced... you mean that the overall effect of all the forces cancels out to have zero net effect

OpenStudy (anonymous):

ohh i see:)

OpenStudy (danjs):



OpenStudy (danjs):

For A) if all the forces have the same magnitude, That doesnt mean the forces are balanced all the time, They would have to have the same magnitude AND the opposite direction (for a two force system)

OpenStudy (anonymous):

I agree with @DanJS . Balanced forces means that the net force is equal to zero. Think of a teeter totter or scale. to balance the scale you need equal force on each side

OpenStudy (danjs):

and distance in that case,, that is a moment or torque

OpenStudy (anonymous):

ahh awesome!! thank you all!! :)

OpenStudy (danjs):

C)The sum of the horizontal forces is zero. Also need sum of the Vertical components to be zero also

OpenStudy (danjs):

D) Pairs in equal magnitude and direction.... that would be additive and not cancel out

OpenStudy (anonymous):

ohh okay, that makes sense now:) thank you very much!! :D

OpenStudy (danjs):

Equilibrium \[\sum F = \sum F _{x} + \sum F _{y} = 0\]

OpenStudy (danjs):

and sum of moments or torque = 0 , but that is a step further

OpenStudy (anonymous):

ahh okay awesome:)

OpenStudy (danjs):

So you decompose your forces using trig rules, into the vertical and horizontal components, then sum them up

OpenStudy (anonymous):

okay yay! thank you:)